OpenStudy (anonymous):

is this right?? x^-1/2 -1/x^1/2 all over 3x^-1/3 -3/x^1/3 =1/x^1/2 -1/x^1/2 all over 3/x^1/3 - 3/x^1/3 =O? thanks guys :) owe you a lot

OpenStudy (anonymous):

yep! As in:\[x^{-a} = \frac{ 1 }{ x^{a} }\]

OpenStudy (anonymous):

@curiousshubham \[x- _{2}^{1} -1/x _{2}^{1} \over 3x- _{3}^{1} -3/x _{3}^{1}\]thankyou

OpenStudy (anonymous):

you are welcomed @xtel .

OpenStudy (anonymous):

@curiousshubham theres the question ^^

hartnn (hartnn):

is the numerator this : \(\huge x^{-1/2}-\frac{1}{x^{1/2}}\) or this : \(\huge \frac{x^{-1/2}-{1}}{x^{1/2}}\)

OpenStudy (anonymous):

If its first one then the solution is indeterminate.

hartnn (hartnn):

yup.

OpenStudy (anonymous):

@halcy0n @curiousshubham yes the first one. it could be turn into this to make it positive? \[1/x _{2}^{1} -1/x _{2}^{1}\]

hartnn (hartnn):

yes, so u get 0 in numerator. but similarly, then u also get 0 in denominator so u have 0/0 which is an indeterminate form

OpenStudy (anonymous):

@hartnn the first one :)

OpenStudy (anonymous):

@xtel So the solution is indeterminate.

OpenStudy (anonymous):

@curiousshubham @hartnn thankyou again guys :)) i owe you a lot